Russell Corporation sold a parcel of land valued at $425,000. Its basis in the land was $359,125. For the land, Russell received $61,500 in cash in year 0 and a note providing that Russell will receive $179,000 in year 1 and $184,500 in year 2 from the buyer (plus reasonable interest on the note).

What is Russell’s recognized gain in year 0, year 1, and year 2?
Answer #1
sale price       425,000.00
basis     (359,125.00)
profit         65,875.00
profit %16%
year 0  61,500*16% =      9,840.00
year 1 179,000*16%=    28,640.00
year 2 185,500*16%=    29,520.00


answered by: olessia silakova